# Disen Lan    Weigao Sun    Jiaxi Hu    Jusen Du    Yu Cheng

### v1

@@ Appendix C Full Results
LoLCATs-Llama-3-8B 80.1 80.4 53.5 63.4 72.9 23.0 62.2 70.0
Liger-GLA-Llama-3-8B (Ours) 80.0 78.7 51.9 76.6 71.4 43.4 67.0 71.7
@@
Model PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Liger-GLA 80.0 78.7 51.9 76.6 71.4 43.4 67.0 71.7
Liger-HGRN2 79.2 76.8 48.5 74.4 68.8 36.2 64.0 69.5
Liger-GSA 79.5 78.5 49.4 74.5 70.5 41.2 65.6 70.5
Table 9: Full Results on Different Gated Linear Recurrent Model Variants with Liger. Liger can be applied to the efficient linearization of various linear recurrent structures with gating mechanism and achive high quality performance recovery.
@@
Table 10: Full Results on Ablation Study. We linearize Llama-3-8B into Gated Linear Attention (GLA) to evaluate the key components of Liger. We report Validation perplexity (PPL.) on cleaned alpaca dataset after Liger linearization and the average performance on language modeling and under-standing tasks.

### v2

@@ Appendix E Full Results
Liger-GLA 80.3 81.1 52.5 76.3 72.0 43.4 67.6 72.4
Liger-HGRN2 79.2 76.8 48.5 74.4 68.8 36.2 64.0 69.5
Liger-GSA 79.5 78.5 49.4 74.5 70.5 41.2 65.6 70.5
Table 10: Full Results on Different Gated Linear Recurrent Model Variants with Liger. Liger can be applied to the efficient linearization of various linear recurrent structures with gating mechanism and achieve high quality performance recovery.
Model PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
@@
LoLCATs-Llama-3-8B 80.1 80.4 53.5 63.4 72.9 42.8 65.5 70.0
Liger-GLA-Llama-3-8B (Ours) 80.3 81.1 52.5 76.3 72.0 43.4 67.6 72.4
@@
- w/o GLA 3.01 79.8 80.5 52.4 76.3 72.4 37.0 66.2 72.0
Table 12: Full Results on Ablation Study. We linearize Llama-3-8B to Gated Linear Attention (GLA) to evaluate the key components of Liger. We report Validation perplexity (PPL.) on cleaned alpaca dataset after Liger linearization and the average performance on language modeling and under-standing tasks.